Again I beleive I have a, b done but need help with c, can someone help please?
2. As you are leaving the community, you notice another new community just down the street. Because you are in the area, you decide to inquire about it.
a) The sales representative here tells you they also have two floor plans available, but they only have 38 homes available. Write an equation that illustrates the situation. Use x and y to denote floor plan #1 and floor plan #2 respectively. x + y = 38
b) The representative tells you that floor plan #1 sells for $175,000 and floor plan #2 sells for $200,000. She also mentions that all the available houses combined are worth $7,200,000. Write an equation that illustrates this situation. Use the same variables you used in part a. $175000x + $200000y = 7200000
c) Use elimination to determine how many houses with each floor plan are available. Explain how you arrived at your answer.
Answer by stanbon(75887) 
 x + y = 38
$175000x + $200000y = 7200000
c)Use elimination to determine how many houses with each floor plan are available. Explain how you arrived at your answer.
---------------
Simplify the 2nd equation by dividing thru by 5000 to get:
35x + 40y = 1440
x + y = 38
-------------------
Multiply thru the 2nd equation by 35 to get:
35x + 35y = 35*38
35x + 40y = 1440
--------------------
Subtract the 1st from the 2nd equation to get:
5y = 10
y = 2 (# of floor plan #1 houses)
----------
Substitute into x + y = 38 to solve for "x":
x + 2 = 38
x = 36 (# of floor plan #2 houses)
======================
Cheers,
Stan H.
| |
|